OpenStudy (anonymous):

On the way to the moon, the Apollo astro- nauts reach a point where the Moon’s gravi- tational pull is stronger than that of Earth’s. Find the distance of this point from the center of the Earth. The masses of the Earth and the Moon are 5.98 × 1024 kg and 7.36 × 1022 kg, respectively, and the distance from the Earth to the Moon is 3.84 × 108 m. Answer in units of m

OpenStudy (john_es):

You must find the point of equilibrium between the two forces, \[G\frac{M_Tm_s}{(R-x)^2}=G\frac{M_Lm_s}{x^2}\] \[\frac{M_T}{(R-x)^2}=\frac{M_L}{x^2}\]So, \[x=R\frac{\sqrt{M_LM_T}-M_L}{M_T-M_L}\]And R is the distance between Earth and Moon.

OpenStudy (john_es):

The result should be, \[x=3.83\cdot10^7\ m\]from the center of the Moon, and \[R-x=3.46\cdot10^8\ m\] from the center of the Earth.

OpenStudy (anonymous):

What would the acceleration of the astro- naut be due to the Earth’s gravity at this point if the moon was not there? The value of the universal gravitational constant is 6.672 × 10−11 N · m2/kg2. Answer in units of m/s2

OpenStudy (john_es):

As the distance from the center of the Earth is the number we found before, \[d=R-x=3.46\cdot10^8\ m\]The acceleration at this point is, \[g=G\frac{M_T}{d^2}=3.33\cdot10^{-3}\ m/s^2\]

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (john_es):

You're welcome.